# Modeling Nonlinear Oscillator Networks Using Physics-Informed Hybrid Reservoir Computing

### v1

@@ Discussion
While the performance advantage of hybrid RC over standard RC and the base ODE is known from previous work [29], we have extended upon this in two ways. First, we recapitulated this result but applied to the important example of an oscillating system. Second, we introduced a new task, the residual physics task, to mimic the situation where the interaction between oscillators is not fully known. This was found to be a more stringent test than the parameter error task, however we still observe higher performance from the hybrid RC than the standard RC or base ODE. As hybrid RCs retain the small size and low training complexity of standard RCs, these results suggest they are a promising architecture for efficient surrogate modeling of NLONs.
@@
Notably, although the hybrid RC had only low valid times on the heteroclinic cycles and partial synchrony regimes, it consistently outperformed the standard RC. This suggests that the hybrid RC does capture some aspects of the residual physics, with, for example, a relative performance gain of $45.8\%$ on the heteroclinic cycles regime. Indeed, although the valid times are low for these regimes, they may be sufficient for control applications with short time horizon requirements.
@@
It is interesting to note the drastic difference in predictions made by the base ODE model and the hybrid RC in the out-of-domain regimes in the residual physics task (Supplementary Fig. S12 and S13). The ODE is in a synchronous regime, and thus predicts synchronized, slow trajectories. This is completely at odds with the ground truth for the heteroclinic cycles or partial synchrony regimes. The addition of the reservoir in the hybrid RC changes this, giving oscillations that much better capture the gross features of the ground truth behavior. In the heteroclinic cycles regime, the hybrid RC more accurately recreates the underlying frequency of the main cluster, and produces oscillators that briefly leave this attracting group. In the partial synchrony regime, the hybrid RC again more closely reproduces the underlying oscillation frequency of the distribution. The hybrid RC does struggle to capture some complex features of the dynamics associated with these regimes, such as the timing of oscillator departure in the heteroclinic cycles regime, and the partially-synchronized, distributional behavior of the partial synchrony regime. When they fail, both the standard and hybrid RC can produce noisy, high-frequency oscillations, or steady state trajectories (Supplementary Fig. S15). As such, the robustness of the hybrid RC may be useful in safety critical applications, where failure through decay or growth of oscillations could be catastrophic.
@@
Over-fitting is a risk for any data-based modeling but is a particular hazard when using models for control. In particular, the standard RC requires careful tuning of the regularization strength to avoid this whereas the hybrid RC does not, even when using an expert model with substantial missing components. In particular, this occurs on the heteroclinic cycles and partial synchrony regimes, where the performance of the hybrid RC is good across a broader range of regularization strengths, above and below the optimum value required for the standard RC. The output matrix of the hybrid RC balances attention between the reservoir activity and the expert model’s predictions. By providing extra states for weight assignment containing some correct features of the ground truth, the hybrid RC reduces the influence of the training data and the risk of over-fitting.
@@
We have shown that the scaling of the reservoir weights has a strong effect on prediction quality for the standard RC; the hybrid RC is also affected, but in a different and less pronounced way. The spectral radius, controlling the scale of the internal connectivity matrix, is particularly influential. This is a well documented phenomenon for a standard RC [36, 37]. The scaling parameters are tied to prediction quality through their impact on the RC’s Lyapunov exponents, and conditional Lyapunov exponents (CLEs). These are features of dynamical systems that determine the structure of phase space, and any constituent attractors. Primarily, they are critical for the satisfaction of the echo state property, a fundamental property for the function of an RC that imbues a reservoir with a fading memory and capacity for generalized signal-induced synchronization [25, 38]. Heuristically it has been suggested that a spectral radius below one is required to have the echo state property, however this is neither necessary or sufficient [33]. The spectral radius is however correlated with the magnitude of the RC’s CLEs [39], and we observe failure of the standard RC for spectral radii greater than one in accordance with this. The hybrid RC does not respond in the same way. Lyapunov exponents and CLEs are also key for successful attractor reconstruction [40, 39]. Although our performance metric is reporting short-term prediction accuracy and not the long-term, ergodic property accuracy, the attractor climate, we suggest that the hybrid RC’s divergent response to increasing spectral radius could be a result of a different correspondence between spectral radius and the CLEs leading to successful attractor reconstruction where the standard RC has failed. Future work should aim to identify the mechanism with which this different behavior occurs.
@@
The scaling parameters have also been implicated in the connection between RC and delay embedding [41, 31]. The high performance of both standard and hybrid RCs at low spectral radius could suggest that they do not require a long memory for the tasks considered here. This is supported by the fact that the ground truth systems we are trying to predict are 1st order ODEs. To investigate whether the reservoirs might benefit from richer representations of fewer past states instead of long memory, we tested reservoirs that completed multiple internal updates per step. These “multi-step” reservoirs are similar to the “drifting-state” reservoirs in  [42] except only the final internal state is used for output computation; after multiple non-linear updates have been applied forming a more complex representation. We found no consistent effect on performance across various dynamical system forecasting tasks (Supplementary Fig. S16). The extra internal steps may produce an effective spectral radius that can become too large or too small, contingent on the original spectral radius setting. This could counteract any benefit of having multiple internal steps. It is also possible that the representations produced by a single step of the reservoir are already rich enough for the dynamical systems considered.
The failure of both the standard and hybrid RC to predict asynchronous dynamics in our study is most likely due to the slow dynamics of that regime (Supplementary Fig. S17). This is not evident in the other three regimes, where the fast dynamics quickly results in the separation of points, and exploration of phase space. The slow movement of the asynchronous dynamics creates a training span in a distinct region of phase space to the test spans. Across both the standard and hybrid RC, this appears to be catastrophic for learning, and subsequent performance. Extending the training span, or training on disjoint training spans that better explore phase space are possible solutions.
@@
The improved performance of the hybrid RC over the standard RC of course comes at the cost of increased computational complexity. This is due to the additional ODE integrator steps being carried out during training and prediction. Further work must evaluate, on a case-by-case basis, whether the performance improvement justifies the cost, and whether the computational cost alone rules out chosen control applications.
@@
All code is available under an MIT License on Github (https://github.com/AJS50/Hybrid_RC_for_NLONS_paper_code).
@@ Acknowledgments
Supplementary Figure S11: Example asynchronous regime phase-component trajectories forecast by each model compared to the ground truth from the residual physics task’s bi-harmonic Kuramoto model. Base ODE (black), standard RC (blue), hybrid RC (red), ground truth (dashed purple). The slow, synchronous base ODE model forecast achieves a higher valid time than the hybrid and standard RC but is clearly failing to capture the main feature of the dynamics once this time has elapsed. The standard RC also fails, but through what appears to be falling into a fixed point solution. Parameters from the first index of the input scaling sweep corresponding to an input scaling of $0.05$, all other parameters at baseline.
@@
Supplementary Figure S16: Multi-step reservoir investigation. Mean valid times achieved by 30 multi-step reservoir instantiations as the number of internal steps varies on the Lorenz system (spectral radius $0.5$, input scaling $0.15$, regularization $0.000001$, others at baseline)(blue), and the heteroclinic cycles regime under different parameter tunings. For un-tuned parameters (spectral radius $0.5$, input scaling $0.15$, regularization $0.000001$, others at baseline)(red), seven internal steps is a local optimum, significantly more than the usual single step. When the parameters are tuned for one internal step however (spectral radius $0.05$, input scaling $0.05$, regularization $0.0001$) (green) this optimum is only two steps. Subsequently, the performance drops off rapidly as the number of internal steps is increased. This may be due to the spectral radius being smaller in the optimized case than the untuned case such that the extra internal steps form an effective spectral radius that is too small. When the spectral radius is adjusted from the tuned value such that the effective spectral radius, defined as the spectral radius of $\textbf{A}^{m}$ where $\mathbf{A}$ is the internal reservoir connectivity matrix, and $m$ is the number of internal steps, is kept constant (at 0.05), the rapid decrease in valid time is ameliorated (purple). Maximum valid times achieved across these cases for the heteroclinic cycles regime do not differ greatly, suggesting optimization of the spectral radius (and other parameters, as conducted in this study) may be equivalent to optimizing the number of internal steps. Dots: individual reservoir instantiation valid time on single test span. Solid/Dashed lines: mean valid time across reservoirs. Shaded regions: one standard deviation across reservoirs.
@@
Supplementary Figure S17: Space-time separation plots of the ground truth trajectories of each dynamical regime of the bi-harmonic Kuramoto model from the residual physics task. These plot the pairwise separation between all pairs of points in the trajectory, with Euclidean distance on the y-axis and time on the x-axis. The asynchronous regime has a protracted region where for small time intervals, the Euclidean distance between points is close. This demonstrates how slow the trajectories are in comparison to the other regimes. We suggest that this is the cause of the training and prediction failures of both the standard and hybrid RCs on the asynchronous regime.

### v2

@@ Discussion
[…]
@@
Notably, although both the standard and hybrid RC had low valid times on the heteroclinic cycles and partial synchrony regimes, the hybrid RC consistently outperformed the standard RC. This suggests that the hybrid RC captures some aspects of the residual physics, with, for example, a relative maximum performance gain of $45.8\%$ on the heteroclinic cycles regime. Although the valid times are low for these regimes, they may be sufficient for control applications with short time horizon requirements.
@@
There are clear limitations of the valid time metric for any application requiring long term, qualitative prediction accuracy, such as attractor reconstruction tasks. This is evident in the qualitative differences between the base ODE and the hybrid RC’s trajectory predictions for asynchronous, heteroclinic cycles, and partial synchrony regimes (Supplementary Fig. S11, S12, and S13), which do not align with their relative performance with valid time. The ODE quickly falls into a synchronous regime, predicting synchronized, slow trajectories at odds with the ground truth for the asynchronous, heteroclinic cycles, or partial synchrony regimes. The addition of the reservoir in the hybrid RC changes this, giving oscillations that much better capture the gross features of the ground truth behavior. In the heteroclinic cycles regime, the hybrid RC more accurately recreates the underlying frequency of the main cluster, and produces oscillators that briefly leave this attracting group. In the partial synchrony regime, the hybrid RC again more closely reproduces the underlying oscillation frequency of the distribution. However it struggles to capture some complex features of these regimes, such as oscillator departure timing in the heteroclinic cycles regime, and the distributional behavior of the partial synchrony regime. When they fail, both the standard and hybrid RC can produce noisy, high-frequency oscillations, or steady state trajectories (Supplementary Fig. S16). As such, the robustness of the hybrid RC may be useful in safety critical applications, where oscillation decay or growth could be catastrophic.
@@
Over-fitting is a risk for any data-based modeling but is a particular hazard when using models for control. The standard RC requires careful tuning of the regularization strength to avoid this whereas the hybrid RC does not, even when using an expert model missing substantial components. This is particularly true for the heteroclinic cycles and partial synchrony regimes, where the performance of the hybrid RC is good across a broader range of regularization strengths, above and below the optimum value required for the standard RC. The output matrix of the hybrid RC balances attention between the reservoir activity and the expert model’s predictions. By providing extra states for weight assignment containing some correct features of the ground truth, the hybrid RC reduces training data influence and over-fitting risk. This is a delicate, task dependent, balance, as evidenced by Figure 9 (b, f, j, n), where the standard RC outperforms the hybrid RC, but the hybrid RC performance increases with increasing regularization strength. Inspection of the predicted trajectories for this regime (Supplementary Fig. S11), indicates the hybrid RC may be biased too far towards the synchronous behavior of its expert model with the baseline regularization setting.
@@
We have shown that the scaling of the reservoir weights has a strong effect on prediction quality for the standard RC; the hybrid RC is also affected, but in a different and less pronounced way. The spectral radius, controlling the scale of the internal connectivity matrix, is particularly influential. This is a well-documented phenomenon for a standard RC [39, 40]. The scaling parameters are tied to prediction quality through their impact on the RC’s Lyapunov exponents, and conditional Lyapunov exponents (CLEs). These are features of dynamical systems that determine the structure of phase space, and any constituent attractors. Primarily, they are critical for the satisfaction of the echo state property, a fundamental property for the function of an RC that imbues a reservoir with a fading memory and capacity for generalized signal-induced synchronization [25, 41]. Heuristically, it has been suggested that a spectral radius below one is required to have the echo state property, however this is neither necessary nor sufficient [35]. The spectral radius is however correlated with the magnitude of the RC’s CLEs [42], and we observe failure of the standard RC for spectral radii greater than one in accordance with this.
The hybrid RC does not respond in the same way. Lyapunov exponents and CLEs are also key for successful attractor reconstruction [43, 42]. Although our performance metric is reporting short-term prediction accuracy and not the long-term ergodic property accuracy, the attractor climate, we suggest that the hybrid RC’s divergent response to increasing spectral radius could be a result of a different correspondence between spectral radius and the CLEs leading to successful attractor reconstruction where the standard RC has failed.
For optimum performance, a spectral radius of zero, corresponding to no memory of past states and no coupling between reservoir nodes, is most effective. This is unsurprising as the ground truth system is first order, and all state information is available to the reservoirs; all the information required for the next step can be obtained from the current step alone.
@@
The scaling parameters have also been implicated in the connection between RC and delay embedding [44, 31]. The high performance of both standard and hybrid RCs at low spectral radius suggests that they do not require a long memory for the tasks considered here. To investigate whether the reservoirs might benefit from richer representations of fewer past states instead of long memory, we tested reservoirs that completed multiple internal updates per step. These “multi-step” reservoirs are similar to the “drifting-state” reservoirs in  [45] except only the final internal state is used for output computation, after multiple non-linear updates have been applied to form a more complex representation. We found no consistent effect on performance across various dynamical system forecasting tasks (Supplementary Fig. S17).
The behavior of the standard and hybrid RC in response to parameter changes is similar across the three non-synchronous regimes. A further test conducted with a slower version of the asynchronous regime resulted in near-zero valid times for both models across all parameters (Supplementary Fig. S18). We believe this was due to the slow, highly correlated, dynamics (Supplementary Fig. S15) creating a training span in a distinct region of phase space from the test spans (Supplementary Fig. S19).
@@
The improved performance of the hybrid RC over the standard RC comes at the cost of increased computational complexity. Further work must evaluate, on a case-by-case basis, whether the performance improvement justifies the cost, and whether the computational cost alone rules out chosen control applications. Extending the hybrid RC architecture to include recent developments in reservoir computing [46, 47, 48] may help to reduce this cost.
@@
The code used to generate and process the results of this study are available from the corresponding author upon reasonable request.
@@ Acknowledgments
Modeling Nonlinear Oscillator Networks Using Physics-Informed Hybrid Reservoir Computing
Andrew Shannon, Conor Houghton, David Barton, Martin Homer
@@
Supplementary Figure S11: Example asynchronous regime phase-component trajectories forecast by each model compared to the ground truth from the residual physics task’s bi-harmonic Kuramoto model. Base ODE (black), standard RC (blue), hybrid RC (red), ground truth (dashed purple). The base ODE model forecast achieves a higher valid time than the hybrid and standard RC but is clearly failing to capture the main feature of the dynamics once this time has elapsed as it synchronizes. The reservoirs can produce more realistic oscillations, with the standard reservoir better capturing the qualitative dynamics for this parameter set. Parameters from the third index of the spectral radius sweep corresponding to a spectral radius of $0.3$, all other parameters at baseline.
@@
Supplementary Figure S15: Example slow asynchronous regime phase-component trajectories forecast by each model compared to the ground truth from the residual physics task’s bi-harmonic Kuramoto model. Base ODE (black), standard RC (blue), hybrid RC (red), ground truth (dashed purple). The slow, synchronous base ODE model forecast achieves a higher valid time than the hybrid and standard RC but is clearly failing to capture the main feature of the dynamics once this time has elapsed. The standard RC also fails, but through what appears to be falling into a fixed point solution. Parameters from the first index of the input scaling sweep corresponding to an input scaling of $0.05$, all other parameters at baseline.
@@
Supplementary Figure S17: Multi-step reservoir investigation. Mean valid times achieved by 30 multi-step reservoir instantiations as the number of internal steps varies on the Lorenz system (spectral radius $0.5$, input scaling $0.15$, regularization $0.000001$, others at baseline)(blue), and the heteroclinic cycles regime under different parameter tunings. For un-tuned parameters (spectral radius $0.5$, input scaling $0.15$, regularization $0.000001$, others at baseline)(red), seven internal steps is a local optimum, significantly more than the usual single step. When the parameters are tuned for one internal step however (spectral radius $0.05$, input scaling $0.05$, regularization $0.0001$) (green) this optimum is only two steps. Subsequently, the performance drops off rapidly as the number of internal steps is increased. This may be due to the spectral radius being smaller in the optimized case than the untuned case such that the extra internal steps form an effective spectral radius that is too small. It is also possible that the representations produced by a single step of the reservoir are already rich enough for the dynamical systems considered. When the spectral radius is adjusted from the tuned value such that the effective spectral radius, defined as the spectral radius of $\textbf{A}^{m}$ where $\mathbf{A}$ is the internal reservoir connectivity matrix, and $m$ is the number of internal steps, is kept constant (at $0.05$), the rapid decrease in valid time is ameliorated (purple). Maximum valid times achieved across these cases for the heteroclinic cycles regime do not differ greatly, suggesting optimization of the spectral radius (and other parameters, as conducted in this study) may be equivalent to optimizing the number of internal steps. Dots: individual reservoir instantiation valid time on single test span. Solid/Dashed lines: mean valid time across reservoirs. Shaded regions: one standard deviation across reservoirs.
Supplementary Figure S18: Residual Physics task parameter sweeps evaluating the hybrid RC’s prediction of NLON trajectories with missing dynamics in its expert model in a slow asynchronous regime. Mean valid time of the prediction of the hybrid RC (red), standard RC (blue), and the base ODE model (black) as four different parameters are varied. Left to right, top to bottom - parameter varied: Spectral radius (a), Input scaling (b), Regularization (c), Reservoir size $D_{r}$ (d). Individual dots are individual reservoir/ODE instantiations (40), each representing the mean NMSE across 60 forecasts, (20 for each realization of a ground truth regime). Solid lines are the mean NMSE across the reservoir/ODE instantiations. Shaded regions are one standard deviation across reservoir/ODE instantiations.
@@
Supplementary Figure S19: Space-time separation plots of the ground truth trajectories of each dynamical regime of the bi-harmonic Kuramoto model from the residual physics task, with the slow asynchronous regime. These plot the pairwise separation between all pairs of points in the trajectory, with Euclidean distance on the y-axis and time on the x-axis. The slow asynchronous regime has a protracted region where for small time intervals, the Euclidean distance between points is close. This demonstrates how slow the trajectories are in comparison to the other regimes. We suggest that this is the cause of the training and prediction failures of both the standard and hybrid RCs on the slow asynchronous regime.
Supplementary Figure S20: Residual Physics task parameter sweeps evaluating the hybrid RC’s prediction of NLON trajectories with missing dynamics in its expert model as the spectral radius of the reservoir is varied down to 0.0. Mean valid time of the prediction of the hybrid RC (red), standard RC (blue), and the base ODE model (black) across the four dynamical regimes. Left to right, top to bottom - dynamical regime: Synchronous (a), Asynchronous (b), Heteroclinic Cycles (c), Partial Synchrony (d). Individual dots are individual reservoir/ODE instantiations (40), each representing the mean NMSE across 60 forecasts, (20 for each realization of a ground truth regime). Solid lines are the mean NMSE across the reservoir/ODE instantiations. Shaded regions are one standard deviation across reservoir/ODE instantiations.
Supplementary Figure S21: Varying the threshold used in the valid time metric computation. Each subplot is at a particular parameter setting from the sweeps in Figure 9, for a given regime. Solid lines: mean mean valid time across tests and instantiations. Dots - mean valid time across tests per instantiation. Shaded region - one standard deviation over instantiations. A threshold of 2.0 is the maximum possible normalized mean square error, occurring when the ground truth state and model prediction are at opposite ends of the state space (20 dimensions). “Accurate” trajectories occur well below this, but this is subjective. A threshold of 0.4 is used in the paper, small variations about this do not significantly affect the ranking of the methods.
Code Implementation
Random Seeds The random number generator used to instantiate each reservoir and ODE parameter settings, as well as the parameter error sampling was of Mersenne Twister type, with a seed set to $1234+parameter\_index$, where the $parameter\_index$ ranged from $1$ to $20$ where the parameter sweep was across 20 points.